# Cutting Through Privacy: A Hyperplane-Based Data Reconstruction Attack in Federated Learning

### v1

@@ 1 Introduction
Federated Learning (FL) enables collaborative machine learning model training across distributed clients without directly sharing data with a central server [McMahan et al., 2017]. Although not sharing data ostensibly guarantees privacy, several attacks have revealed existing privacy vulnerabilities in FL. By inspecting the updates received and using publicly available information, the central server can infer training data membership [Melis et al., 2019, Nasr et al., 2019, Choquette-Choo et al., 2021] or infer clients’ sensitive attributes [Chen et al., 2024, Diana et al., 2024]. These attacks only expose limited information, whereas more advanced data reconstruction attacks enable the server to fully recover private training data points. These attacks have been proposed in two primary settings: the honest-but-curious setting, where the server passively observes updates [Zhu et al., 2019, Zhao et al., 2020, Geiping et al., 2020, Yin et al., 2021, Kariyappa et al., 2023, Dimitrov et al., 2022, 2024], and the malicious setting [Fowl et al., 2022, Boenisch et al., 2023, Zhang et al., 2023, Zhao et al., 2024, Wang et al., 2024], where the server actively manipulates model parameters. In the honest-but-curious setting, existing methods struggle to perfectly recover the training data for batch sizes of a few dozen data points [Dimitrov et al., 2024]) or produce very noisy reconstructions for a large portion of data points in batches exceeding 100 data points [Kariyappa et al., 2023, Fig. 6] and Garov et al. [2024]. In the malicious setting, existing attacks [Wen et al., 2022, Fowl et al., 2022, Boenisch et al., 2023, Zhang et al., 2023] perform better but have not succeeded in perfectly recovering a batch of 256 data points.
@@ 2 Background
where $\mathcal{D}^{u}=\{({\mathbf{x}}_{i}^{u},y_{i}^{u})\}_{i=1}^{n^{u}}$ is the local dataset of client $u$, ${\mathcal{L}^{u}}$ measures the loss of model $\theta$ on points in $\mathcal{D}^{u}$, and $p^{u}$ is a positive weight quantifying the relative importance of client $u$, s.t. $\sum_{u\in\mathcal{U}}p^{u}=1$. Typical choices in FL are $p_{u}=1/|\mathcal{U}|$ and $p_{u}\propto n_{u}$.
@@
While FL preserves data locality, the transmitted updates—whether gradients or model parameters—can leak sensitive information about the clients’ datasets [Melis et al., 2019, Diana et al., 2024, Fowl et al., 2022]. This vulnerability exposes FL to adversarial attacks aimed at reconstructing private data [Zhu et al., 2019, Fowl et al., 2022, Boenisch et al., 2023, Dimitrov et al., 2024]. The success rate of such attacks depends on the adversary’s capabilities.
@@ 2.1 Optimization-based attacks
Early works [Zhu et al., 2019, Zhao et al., 2020, Geiping et al., 2020, Yin et al., 2021, Dimitrov et al., 2022] proposed the recovery of image inputs within the honest-but-curious setting, but were successful only on small ($\leq 50$) batch sizes and, even then, produced noisy reconstructions. Kariyappa et al. [2023] reformulated the reconstruction task as a blind source separation problem, enabling recovery in presence of secure aggregation [Bonawitz et al., 2017], however, they still only obtain very noisy images on batches larger than 200 data points (see Figure 6 in their work).
Garov et al. [2024] tackled the same problem by training an encoder-decoder network, though requiring auxiliary data. While their approach produces high-quality images, it successfully reconstructs only half of the images in a batch of 512. Vero et al. [2023] introduced the first optimization-based attack targeting tabular data and reported a 70% reconstruction success rate on a batch of 128 samples.
Overall, these attacks struggle with the challenge posed by gradient aggregation, and do not scale effectively to large batch sizes without requiring auxiliary data. To address these shortcomings, alternative approaches consider a scenario where the server is malicious. In this context, Wen et al. [2022] proposed a fishing strategy that alters the weights of the classification layer of a neural network to amplify the gradient contribution of a single input within a batch, increasing the success rate of the optimization step. However, their approach relies on estimating the distribution of one input feature, which the authors assume to follow a normal distribution. While this assumption may hold true for image data, it is often unrealistic for tabular datasets, where features can be binary or categorical.
@@ 2.2 Analytical attacks
For simplicity, consider a client training a model where the first layer is a fully connected (FC) layer with input vector ${\mathbf{x}}\in\mathbb{R}^{d}$ and output vector ${\mathbf{z}}\in\mathbb{R}^{N}$. Let $\mathbf{W}\in\mathbb{R}^{N\times d}$ and ${\mathbf{b}}\in\mathbb{R}^{N}$ denote the weight matrix and the bias vector, respectively. Let the output of the $i$-th output neuron in the layer be $z_{i}=\textrm{ReLU}(\mathbf{W}_{i}{\mathbf{x}}+b_{i})$. Given a sample $({\mathbf{x}}_{j},y_{j})$, let $\mathcal{L}_{j}$ and $z_{i,j}$ denote the loss and the output of neuron $i$ for this sample, respectively. The gradient of the loss with respect to $\mathbf{W}_{i}$ and $b_{i}$ can be computed as follows:111 The expression $\frac{\partial\mathbf{f}}{\partial{\mathbf{x}}}$, where $\mathbf{f}:\mathbb{R}^{d}\to\mathbb{R}^{N}$, can also be interpreted as the Jacobian of the function $\mathbf{f}$, with the (less common) convention that $\left(\frac{\partial\mathbf{f}}{\partial{\mathbf{x}}}\right)_{i,j}=\frac{\partial f_{j}}{\partial x_{i}}$. $\frac{\partial\mathcal{L}_{j}}{\partial\mathbf{W}_{i}}=\frac{\partial\mathcal{L}_{j}}{\partial z_{i}}\frac{\partial z_{i}}{\partial\mathbf{W}_{i}}=\frac{\partial\mathcal{L}_{j}}{\partial z_{i}}{\mathbf{x}}_{j}\mathbbm{1}_{z_{i,j}>0}$ and $\frac{\partial\mathcal{L}_{j}}{\partial b_{i}}=\frac{\partial\mathcal{L}_{j}}{\partial z_{i}}\frac{\partial z_{i}}{\partial b_{i}}=\frac{\partial\mathcal{L}_{j}}{\partial z_{i}}\mathbbm{1}_{z_{i,j}>0}$, from which it follows that $\frac{\partial\mathcal{L}_{j}}{\partial\mathbf{W}_{i}}=\frac{\partial\mathcal{L}_{j}}{\partial b_{i}}{\mathbf{x}}_{j}$. Consequently, as long as $\frac{\partial\mathcal{L}_{j}}{\partial b_{i}}\neq 0$ (which requires that ${\mathbf{x}}_{j}$ activates neuron $i$, i.e., $\mathbf{W}_{i}{\mathbf{x}}_{j}+b_{i}>0$), knowing the gradient of the model loss enables exact reconstruction of the input:
@@
However, in practice, gradients are computed on a batch =$\mathcal{B}=\{({\mathbf{x}}_{j},y_{j})\}_{j=1}^{n}$ of size $n$, rather than on a single input. The attacker can then only observe the gradient of the average loss $\mathcal{L}=\frac{1}{n}\sum_{j=1}^{n}\mathcal{L}_{j}$:
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}_{i}}=\frac{1}{n}\sum_{j=1}^{n}\frac{\partial\mathcal{L}_{j}}{\partial b_{i}}{\mathbf{x}}_{j}\,\$ (3)
$\frac{\partial\mathcal{L}}{\partial b_{i}}=\frac{1}{n}\sum_{j=1}^{n}\frac{\partial\mathcal{L}_{j}}{\partial b_{i}}.\$ (4)
If $\frac{\partial\mathcal{L}}{\partial b_{i}}\neq 0$ (i.e., at least one input activates neuron $i$), then the attacker can reconstruct a combination $\mathbf{g}$ of the inputs from this gradient [Zhang et al., 2023, Theorem 1]):
$\mathbf{g}$ $=\frac{\partial{\mathcal{L}}}{\partial{\mathbf{W}_{i}}}\cdot\left(\frac{\partial{\mathcal{L}}}{\partial{b_{i}}}\right)^{-1}=\sum_{j=1}^{n}\alpha_{j}{\mathbf{x}}_{j},$ (5)
@@ 2.2.1 Sparsity-based attacks
As we will demonstrate in Sec. 3, the effectiveness of the sparsity-based attacks on low-dimensional data, such as tabular datasets, is limited.
@@ 2.2.2 Other analytical attacks
Some analytical attacks can recover input data without inducing gradient sparsity. Fowl et al. [2022] assume knowledge of the distribution of a linear combination of the inputs in the local dataset and use this information to select bias values for all neurons, ensuring that multiple combinations in (3) and (4) differ by a single input, which can then be easily decoded. The first round of our parallel search method presented in Sec. 4.4 is similar to their approach but there are two crucial differences: (i) they require prior knowledge of a distribution, while we dynamically discover the appropriate biases, and (ii) they do not present means to retain relevant information across multiple communication rounds, ruling out a search approach (like ours).
@@ 2.2.3 Extension to other architectures
Until now, we have discussed reconstruction attacks targeting the first layer in the network. These attacks can be directly extended if the attacked layer is preceded by other fully connected (FC) layers.
@@
Lastly, Lu et al. [2022] demonstrated that analytical recovery of images is also feasible in attention-based networks.
@@ 3 Limitations of Existing Sparsity-Based Attacks
The sparsity-based attacks proposed by Boenisch et al. [2023], Zhang et al. [2023] attempt to modify the parameters of the FC layer so that each neuron in the layer is activated only by a single input. Given the batch =$\mathcal{B}=\{({\mathbf{x}}_{1},y_{1})\dots,({\mathbf{x}}_{n},y_{n})\}$, neuron $i$ is activated only from inputs ${\mathbf{x}}_{j}$ if $\mathbf{W}_{i}{\mathbf{x}}_{j}+b_{i}>0$ and $\mathbf{W}_{i}{\mathbf{x}}_{k}+b_{i}<0$ for $k\in\{1,\dots,n\}\setminus\{j\}$. In this section we provide a geometric interpretation of the problem and demonstrate that the accuracy of sparsity-based attacks is fundamentally constrained by the dimensionality of the input they aim to reconstruct.
Consider the set of inputs for which the activation value of neuron $i$ is equal to 00, i.e., $\{{\mathbf{x}}\in\mathbb{R}^{d}\mid\mathbf{W}_{i}{\mathbf{x}}+b_{i}=0\}$. This set is a hyperplane in the Euclidean space $\mathbb{R}^{d}$ orthogonal to the vector $\mathbf{W}_{i}$. The attacks proposed by Boenisch et al. [2023], Zhang et al. [2023] can be interpreted as identifying hyperplanes orthogonal to random directions $\mathbf{W}_{i}$ that separate a single input from the rest of the inputs in the batch, which we denote by $\mathcal{B}_{{\mathbf{x}}}=\{{\mathbf{x}}_{1},{\mathbf{x}}_{2},\dots,{\mathbf{x}}_{n}\}$. Given a gradient computed on the batch $\mathcal{B}$, the maximum number of recoverable points is upper-bounded by the number of points in $\mathcal{B}_{{\mathbf{x}}}$ that are linearly separable from all the others, i.e., the number of vertices on the convex hull of $\mathcal{B}_{{\mathbf{x}}}$. By bounding the expected number of vertices, we derive the following theorem (proof in App. A):
@@
    $O(\log^{(d-1)/2}n)$.
Theorem 3.1 shows that, for a fixed input dimension $d$, the expected fraction of reconstructed samples—i.e., the attack success rate—approaches zero asymptotically as the batch size $n$ increases. This result explains the attack performance degradation observed by Boenisch et al. [2023], Zhang et al. [2023] when larger batch sizes are used. Furthermore, the theorem highlights a fundamental limitation in low-dimensional settings: the smaller the input dimension $d$, the lower the success rate. This suggests that while these attacks may be effective for high-dimensional data, their applicability to lower-dimensional data is significantly constrained. Moreover, we observe that even high-dimensional data often resides on manifolds of lower dimensionality [Narayanan and Mitter, 2010, Pope et al., 2021, Brown et al., 2023], potentially further reducing the success rate of these attacks beyond the predictions of Theorem 3.1.
@@ 4.2 Locate and Reconstruct the Inputs
Consider that the client under attack has dataset $\mathcal{B}$ with corresponding inputs $\mathcal{B}_{{\mathbf{x}}}$. For simplicity, we assume that the neural network consists solely of an FC input layer with ReLU activations (with parameters $\mathbf{W}^{(1)}\in\mathbb{R}^{N\times d}$ and ${\mathbf{b}}^{(1)}\in\mathbb{R}^{N}$) and a FC classification layer (with parameters $\mathbf{W}^{(2)}\in\mathbb{R}^{N\times C}$ and ${\mathbf{b}}^{(2)}\in\mathbb{R}^{C}$). The output of the network can be expressed as ${\mathbf{z}}^{(2)}=\mathbf{W}^{(2)}{\mathbf{z}}^{(1)}+{\mathbf{b}}^{(2)}\in\mathbb{R}^{C}$, where ${\mathbf{z}}^{(1)}\leavevmode\nobreak\ =\textrm{ReLU}(\mathbf{W}^{(1)}{\mathbf{x}}+{\mathbf{b}}^{(1)})\in\mathbb{R}^{N}$. The attack can be extended to other architectures as described in other papers (see Sec. 2.2.3).
@@
For each input $({\mathbf{x}}_{j},y_{j})\in\mathcal{B}$ and values $b^{(1)}_{i}$ tested by the attacker, if neuron $i$ is activated by the input ${\mathbf{x}}_{j}$ for a given value of $b^{(1)}_{i}$, the derivative ${\partial\mathcal{L}_{j}}/{\partial b^{(1)}_{i}}$ remains independent of $b^{(1)}_{i}$ and $\alpha_{j}$ (6) is non-negligible.
According to (5), for a given bias value $b^{(1)}_{i,k}$, the server observes the client’s aggregate gradient $\mathbf{g}_{k}=\sum_{j=1}^{n}\alpha_{j,k}{\mathbf{x}}_{j}$ with a non-negligible $\alpha_{j,k}$ for activated input ${\mathbf{x}}_{j}$ as in (6). Given hyperplanes $\mathbf{W}^{(1)}_{i}{\mathbf{x}}+b^{(1)}_{i,k}=0$ and $\mathbf{W}^{(1)}_{i}{\mathbf{x}}+b^{(1)}_{i,k+1}=0$, the attacker can determine whether they enclose at least one point by comparing $\mathbf{g}_{k}$ and $\mathbf{g}_{k+1}$. Indeed, if there is no input point between the two hyperplanes, the same set of inputs activates neuron $i$ and $\alpha_{i,k}=\alpha_{i,k+1}$ (because of Assumption 4.1), then $\mathbf{g}_{k}=\mathbf{g}_{k+1}$. If $\mathbf{g}_{k}\neq\mathbf{g}_{k+1}$, then the server knows that there is at least one input in-between and, by binary search, can progressively isolate the inputs as far as their projections along the direction $(\mathbf{W}_{i}^{(1)})^{\intercal}$ differ by more than a threshold $\varepsilon$ (see App. D for its configuration rule).
After isolating all the input points, the attacker proceeds with their reconstruction as follows. The attacker can reconstruct the inputs sequentially from ${\mathbf{x}}_{1}$ to ${\mathbf{x}}_{n}$, where we relabel the inputs so that ${\mathbf{x}}_{j}\in\mathcal{B}_{{\mathbf{x}}}$ is the input between the two hyperplanes corresponding to $\hat{b}^{(1)}_{i,j-1}$ and $\hat{b}^{(1)}_{i,j}$. Input ${\mathbf{x}}_{1}$ can be directly reconstructed using (2) with the client’s loss gradient computed for $\hat{b}^{(1)}_{i,1}$. We note that the attacker also knows ${\partial\mathcal{L}_{1}}/{\partial b^{(1)}_{i}}$ for $b^{(1)}=\hat{b}^{(1)}_{i,1}$, i.e., when ${\mathbf{x}}_{1}$ activates neuron $i$. Suppose now that the attacker already reconstructed the first $k$ inputs ${\mathbf{x}}_{1},\dots,{\mathbf{x}}_{k}$ and the corresponding partial derivatives ${\partial\mathcal{L}_{1}}/{\partial b^{(1)}_{i}},\dots,{\partial\mathcal{L}_{k}}/{\partial b^{(1)}_{i}}$ when the corresponding inputs activate neuron $i$. Let $h_{k}$ denote the partial derivative of the loss for $\hat{b}^{(1)}_{i,k}$, i.e., $h_{k}={\partial\mathcal{L}}/{\partial b^{(1)}_{i}}$. Due to (4) and Ass. 4.1, the attacker can reconstruct ${\partial\mathcal{L}_{k+1}}/{\partial b^{(1)}_{i}}$ by
@@
Figure 1: Illustration of our attack strategy, where hyperplanes are positioned to isolate a single image within the strip they define.

### v2

@@ 1 Introduction
Federated Learning (FL) enables collaborative machine learning model training across distributed clients without directly sharing data with a central server [McMahan et al., 2017]. Although not sharing data ostensibly guarantees privacy, several attacks have revealed existing privacy vulnerabilities in FL. By inspecting the updates received and using publicly available information, the central server can infer training data membership [Melis et al., 2019, Nasr et al., 2019, Choquette-Choo et al., 2021] or infer clients’ sensitive attributes [Chen et al., 2024, Diana et al., 2025]. These attacks only expose limited information, whereas more advanced data reconstruction attacks enable the server to fully recover private training data points. These attacks have been proposed in two primary settings: the honest-but-curious setting, where the server passively observes updates [Zhu et al., 2019, Zhao et al., 2020, Geiping et al., 2020, Yin et al., 2021, Kariyappa et al., 2023, Dimitrov et al., 2022, 2024], and the malicious setting [Fowl et al., 2022, Boenisch et al., 2023, Zhang et al., 2023, Zhao et al., 2024, Wang et al., 2024, Garov et al., 2024], where the server actively manipulates model parameters. In the honest-but-curious setting, existing methods struggle to perfectly recover the training data for batch sizes of a few dozen data points [Dimitrov et al., 2024]) or produce very noisy reconstructions for a large portion of data points in batches exceeding 100 data points [Kariyappa et al., 2023, Fig. 6]. In the malicious setting, existing attacks [Wen et al., 2022, Fowl et al., 2022, Boenisch et al., 2023, Zhang et al., 2023, Garov et al., 2024] perform better but have not succeeded in perfectly recovering a batch of 256 data points.
@@ 2 Background
where $\mathcal{D}^{u}=\{({\mathbf{x}}_{i}^{u},y_{i}^{u})\}_{i=1}^{n^{u}}$ is the local dataset of client $u$, $\mathcal{L}^{u}$} measures the loss of model $\theta$ on points in $\mathcal{D}^{u}$, and $p^{u}$ is a positive weight quantifying the relative importance of client $u$, s.t. $\sum_{u\in\mathcal{U}}p^{u}=1$. Typical choices in FL are $p^{u}=1/|\mathcal{U}|$ and $p^{u}\propto n^{u}$.
@@
While FL preserves data locality, the transmitted updates—whether gradients or model parameters—can leak sensitive information about the clients’ datasets [Melis et al., 2019, Diana et al., 2025, Fowl et al., 2022]. This vulnerability exposes FL to adversarial attacks aimed at reconstructing private data [Zhu et al., 2019, Fowl et al., 2022, Boenisch et al., 2023, Dimitrov et al., 2024]. The success rate of such attacks depends on the adversary’s capabilities. In what follows, we consider the attacker to be the server.
@@ 2.1 Optimization-based attacks
Early works [Zhu et al., 2019, Zhao et al., 2020, Geiping et al., 2020, Yin et al., 2021, Dimitrov et al., 2022] proposed the recovery of image inputs within the honest-but-curious setting, but were successful only on small ($\leq 50$) batch sizes and, even then, produced noisy reconstructions. Kariyappa et al. [2023] reformulated the reconstruction task as a blind source separation problem, enabling recovery in presence of secure aggregation [Bonawitz et al., 2017], however, they still only obtained very noisy images on batches larger than 200 data points (see Figure 6 in their work).
Vero et al. [2023] introduced the first optimization-based attack targeting tabular data and reported a 70% reconstruction success rate on a batch of 128 samples.
Overall, these attacks struggle with the challenge posed by gradient aggregation, and do not scale effectively to large batch sizes without requiring auxiliary data. To address these shortcomings, alternative approaches consider a scenario where the server is malicious. In this context, Wen et al. [2022] proposed a fishing strategy that alters the weights of the classification layer of a neural network to amplify the gradient contribution of a single input within a batch, increasing the success rate of the optimization step. However, their approach relies on estimating the distribution of one input feature, which the authors assume to follow a normal distribution. While this assumption may hold true for image data, it is often unrealistic for tabular datasets, where features can be binary or categorical. Alternatively, Garov et al. [2024] designed an encoder-decoder architecture to isolate individual input gradients from the aggregated update in a latent space using an encoder to reduce the detectability, and subsequently reconstruct the original input with a decoder. Nonetheless, their method depends on access to auxiliary data for training the encoder-decoder network and struggles to achieve high-quality recovery for approximately half of the images in a batch of 512 samples.
@@ 2.2 Analytical attacks
For simplicity, consider a client training a model where the first layer is a fully connected (FC) layer with input vector ${\mathbf{x}}\in\mathbb{R}^{d}$ and output vector ${\mathbf{z}}\in\mathbb{R}^{N}$. Let $\mathbf{W}\in\mathbb{R}^{N\times d}$ and ${\mathbf{b}}\in\mathbb{R}^{N}$ denote the weight matrix and the bias vector, respectively. Let the output of the $i$-th neuron in the layer be $z_{i}=\textrm{ReLU}(\mathbf{W}_{i}{\mathbf{x}}+b_{i})$. Given a sample $({\mathbf{x}}_{j},y_{j})$, let $\mathcal{L}_{j}$ and $z_{i,j}$ denote the loss and the output of neuron $i$ for this sample, respectively. The gradient of the loss with respect to $\mathbf{W}_{i}$ and $b_{i}$ can be computed as follows:111 The expression $\frac{\partial\mathbf{f}}{\partial{\mathbf{x}}}$, where $\mathbf{f}:\mathbb{R}^{d}\to\mathbb{R}^{N}$, can also be interpreted as the Jacobian of the function $\mathbf{f}$, with the (less common) convention that $\left(\frac{\partial\mathbf{f}}{\partial{\mathbf{x}}}\right)_{i,j}=\frac{\partial f_{j}}{\partial x_{i}}$. $\frac{\partial\mathcal{L}_{j}}{\partial\mathbf{W}_{i}}=\frac{\partial\mathcal{L}_{j}}{\partial z_{i}}\frac{\partial z_{i}}{\partial\mathbf{W}_{i}}=\frac{\partial\mathcal{L}_{j}}{\partial z_{i}}{\mathbf{x}}_{j}\mathbbm{1}_{z_{i,j}>0}$ and $\frac{\partial\mathcal{L}_{j}}{\partial b_{i}}=\frac{\partial\mathcal{L}_{j}}{\partial z_{i}}\frac{\partial z_{i}}{\partial b_{i}}=\frac{\partial\mathcal{L}_{j}}{\partial z_{i}}\mathbbm{1}_{z_{i,j}>0}$, from which it follows that $\frac{\partial\mathcal{L}_{j}}{\partial\mathbf{W}_{i}}=\frac{\partial\mathcal{L}_{j}}{\partial b_{i}}{\mathbf{x}}_{j}$. Consequently, as long as $\frac{\partial\mathcal{L}_{j}}{\partial b_{i}}\neq 0$ (which requires that ${\mathbf{x}}_{j}$ activates neuron $i$, i.e., $\mathbf{W}_{i}{\mathbf{x}}_{j}+b_{i}>0$), knowing the gradient of the model loss enables exact reconstruction of the input:
@@
However, in practice, gradients are computed on a batch $\mathcal{B}=\{({\mathbf{x}}_{j},y_{j})\}_{j=1}^{n}$ of size $n$, rather than on a single input. The attacker can then only observe the gradient of the average loss $\mathcal{L}=\frac{1}{n}\sum_{j=1}^{n}\mathcal{L}_{j}$:
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}_{i}}=\frac{1}{n}\sum_{j=1}^{n}\frac{\partial\mathcal{L}_{j}}{\partial b_{i}}{\mathbf{x}}_{j}\,\\$ (3)
$\frac{\partial\mathcal{L}}{\partial b_{i}}=\frac{1}{n}\sum_{j=1}^{n}\frac{\partial\mathcal{L}_{j}}{\partial b_{i}}.\\$ (4)
If $\frac{\partial\mathcal{L}}{\partial b_{i}}\neq 0$—implying that at least one input activates neuron ii—the attacker can reconstruct a combination $\mathbf{g}_{i}$ of the inputs from this gradient [Zhang et al., 2023, Theorem 1]:
$\mathbf{g}_{i}$ $=\frac{\partial{\mathcal{L}}}{\partial{\mathbf{W}_{i}}}\cdot\left(\frac{\partial{\mathcal{L}}}{\partial{b_{i}}}\right)^{-1}=\sum_{j=1}^{n}\alpha_{j}{\mathbf{x}}_{j},$ (5)
@@ 2.2.1 Sparsity-Based Attacks
As we will demonstrate in Sec. 3, sparsity-based attacks are significantly less effective on low-dimensional data, such as tabular datasets.
@@ 2.2.2 Other Analytical Attacks
Some analytical attacks can recover input data without inducing gradient sparsity. Fowl et al. [2022] assume prior knowledge of the distribution of a linear combination of client inputs and configure neuron biases such that the gradients with respect to different neurons’ weights differ by only a single input, which can be recovered using (3) and (4). The first round of our parallel search method presented in Sec. 4.4 is similar to their approach but there are two crucial differences: (i) they require prior knowledge of a distribution, while we dynamically discover the appropriate biases, and (ii) they do not propose any iterative procedure that enables a bias search across multiple communication rounds, as we do.
@@ 2.2.3 Extension to other architectures
Until now, we have discussed reconstruction attacks targeting the first layer in the network. These attacks can be directly extended when the attacked layer is preceded by additional FC layers.
@@
Other works have specifically targeted transformer-based models. Lu et al. [2022] demonstrated that analytical image recovery is feasible even in attention-based networks; however, their attack is restricted to a batch size of one. Fowl et al. [2023] and Chu et al. [2023] extended their earlier work [Fowl et al., 2022] to extract textual data from clients; however, their methods do not extend to other data modalities, such as images.
@@ 3 Limitations of Existing Sparsity-Based Attacks
The sparsity-based attacks proposed by Boenisch et al. [2023], Zhang et al. [2023] attempt to modify the parameters of the FC layer so that each neuron in the layer is activated only by a single input. Given the batch $\mathcal{B}=\{({\mathbf{x}}_{1},y_{1})\dots,({\mathbf{x}}_{n},y_{n})\}$, neuron $i$ is activated only from input ${\mathbf{x}}_{j}$ if $\mathbf{W}_{i}{\mathbf{x}}_{j}+b_{i}>0$ and $\mathbf{W}_{i}{\mathbf{x}}_{k}+b_{i}<0$ for $k\in\{1,\dots,n\}\setminus\{j\}$. In this section we provide a geometric interpretation of the problem and demonstrate that the accuracy of sparsity-based attacks is fundamentally constrained by the dimensionality of the input they aim to reconstruct.
Consider the set of inputs for which the activation value of neuron $i$ is equal to 0, i.e., $\{{\mathbf{x}}\in\mathbb{R}^{d}\mid\mathbf{W}_{i}{\mathbf{x}}+b_{i}=0\}$. This set is a hyperplane in the Euclidean space $\mathbb{R}^{d}$ orthogonal to the vector $\mathbf{W}_{i}$. The attacks proposed by Boenisch et al. [2023], Zhang et al. [2023] can be interpreted as identifying hyperplanes orthogonal to random directions $\mathbf{W}_{i}$ that separate a single input from the rest of the inputs in the batch, which we denote by $\mathcal{B}_{{\mathbf{x}}}=\{{\mathbf{x}}_{1},{\mathbf{x}}_{2},\dots,{\mathbf{x}}_{n}\}$. Given a gradient computed on the batch $\mathcal{B}$, the maximum number of recoverable points is upper-bounded by the number of points in $\mathcal{B}_{{\mathbf{x}}}$ that are linearly separable from all the others, i.e., the number of vertices on the convex hull of $\mathcal{B}_{{\mathbf{x}}}$. By bounding the expected number of vertices, we derive the following theorem (proof in App. A):
@@
    $O(\log^{(d-1)/2}n)$,
respectively for the three distributions above.
Theorem 3.1 shows that, for a fixed input dimension $d$, the expected fraction of reconstructed samples—i.e., the attack success rate—approaches zero asymptotically as the batch size $n$ increases. This result explains the attack performance degradation observed by Boenisch et al. [2023], Zhang et al. [2023] when larger batch sizes are used. Furthermore, the theorem highlights a fundamental limitation in low-dimensional settings: the smaller the input dimension $d$, the lower the success rate. This indicates that while such attacks can succeed in high-dimensional regimes, their effectiveness diminishes substantially when applied to low-dimensional data. Moreover, high-dimensional data often lies on manifolds of much lower intrinsic dimensionality [Narayanan and Mitter, 2010, Pope et al., 2021, Brown et al., 2023], potentially reducing the success rate of these attacks even further than predicted by Theorem 3.1.
@@ 4.2 Locate and Reconstruct the Inputs
Consider that the client under attack has dataset $\mathcal{B}$ with corresponding inputs $\mathcal{B}_{{\mathbf{x}}}$. For simplicity, we assume that the neural network consists solely of a FC input layer with ReLU activations (with parameters $\mathbf{W}^{(1)}\in\mathbb{R}^{N\times d}$ and ${\mathbf{b}}^{(1)}\in\mathbb{R}^{N}$) and a FC classification layer (with parameters $\mathbf{W}^{(2)}\in\mathbb{R}^{C\times N}$ and ${\mathbf{b}}^{(2)}\in\mathbb{R}^{C}$). The output of the network can be expressed as ${\mathbf{z}}^{(2)}=\mathbf{W}^{(2)}{\mathbf{z}}^{(1)}+{\mathbf{b}}^{(2)}\in\mathbb{R}^{C}$, where ${\mathbf{z}}^{(1)}\leavevmode\nobreak\ =\textrm{ReLU}(\mathbf{W}^{(1)}{\mathbf{x}}+{\mathbf{b}}^{(1)})\in\mathbb{R}^{N}$. The attack can be extended to other architectures as described in other papers (see discussion in Sec. 2.2.3).
@@
For each input $({\mathbf{x}}_{j},y_{j})\in\mathcal{B}$ and each value of $b^{(1)}_{i}$ tested by the attacker, if neuron $i$ is activated by the input ${\mathbf{x}}_{j}$ for a given value of $b^{(1)}_{i}$, the derivative ${\partial\mathcal{L}_{j}}/{\partial b^{(1)}_{i}}$ remains independent of $b^{(1)}_{i}$.
According to (5), for a given bias value $b^{(1)}_{i,k}$, the server observes the client’s aggregate gradient $\mathbf{g}_{k}=\sum_{j=1}^{n}\alpha_{j,k}{\mathbf{x}}_{j}$ Given hyperplanes $\mathbf{W}^{(1)}_{i}{\mathbf{x}}+b^{(1)}_{i,k}=0$ and $\mathbf{W}^{(1)}_{i}{\mathbf{x}}+b^{(1)}_{i,k+1}=0$, the attacker can determine whether they enclose at least one point by comparing $\mathbf{g}_{k}$ and $\mathbf{g}_{k+1}$. Indeed, if there is no input point between the two hyperplanes, the same set of inputs activates neuron $i$ and $\alpha_{i,k}=\alpha_{i,k+1}$ (because of Assumption 4.1), then $\mathbf{g}_{k}=\mathbf{g}_{k+1}$. If $\mathbf{g}_{k}\neq\mathbf{g}_{k+1}$, then the server knows that there is at least one input in-between and, by binary search, can progressively isolate the inputs as far as their projections along the direction $(\mathbf{W}_{i}^{(1)})^{\intercal}$ differ by more than a threshold $\varepsilon$ (see App. D for its configuration rule).
After isolating all the input points, the attacker can reconstruct the inputs sequentially from ${\mathbf{x}}_{1}$ to ${\mathbf{x}}_{n}$, where we relabel the inputs so that ${\mathbf{x}}_{j}\in\mathcal{B}_{{\mathbf{x}}}$ is the input between the two hyperplanes corresponding to $\hat{b}^{(1)}_{i,j-1}$ and $\hat{b}^{(1)}_{i,j}$. Input ${\mathbf{x}}_{1}$ can be directly reconstructed using (2) with the client’s loss gradient computed for $\hat{b}^{(1)}_{i,1}$. We note that the attacker also knows ${\partial\mathcal{L}_{1}}/{\partial b^{(1)}_{i}}$ for $b^{(1)}=\hat{b}^{(1)}_{i,1}$, i.e., when ${\mathbf{x}}_{1}$ activates neuron $i$. Suppose now that the attacker already reconstructed the first $k$ inputs ${\mathbf{x}}_{1},\dots,{\mathbf{x}}_{k}$ and the corresponding partial derivatives ${\partial\mathcal{L}_{1}}/{\partial b^{(1)}_{i}},\dots,{\partial\mathcal{L}_{k}}/{\partial b^{(1)}_{i}}$ when the corresponding inputs activate neuron $i$. Let $h_{k}$ denote the partial derivative of the loss for $\hat{b}^{(1)}_{i,k}$, i.e., $h_{k}={\partial\mathcal{L}}/{\partial b^{(1)}_{i}}$. Due to (4) and Ass. 4.1, the attacker can reconstruct ${\partial\mathcal{L}_{k+1}}/{\partial b^{(1)}_{i}}$ by
@@
Figure 1: Illustration of our attack strategy, in which hyperplanes are positioned to isolate a single image within the strip they define. The light-blue region represents the half-space where $\mathbf{W}^{(1)}_{i}{\mathbf{x}}+b_{1}^{(1)}>0$.